A. If you want to earn 9% annual simple interest on an investment, how much should you pay for a note that will be worth $24,400.00 in 2 months. You should pay $? B. If $9,000.00 is what you invested at 16% annual simple interest, how long does it take to be worth $37,800.00. It will take years (round to the nearest whole)?

B.If $9,000.00 is invested at 16% annual simple interest, how long does it take to be worth $37,800.00.
It will take  years (round to the nearest whole)?
Answer #1

A.

First, converting R percent to r a decimal
r = R/100 = 9%/100 = 0.09 per year.

Putting time into years for simplicity,
2 months / 12 months/year = 0.166667 years

Solving our equation:
P = 24400 / ( 1 + (0.09 × 0.166667)) = 24039.408156471
P = $24,039.41

B.

First, converting R percent to r a decimal
r = R/100 = 16%/100 = 0.16 per year,
then, solving our equation

t = (1/0.16)((37800/9000) - 1) = 20
t = 20 years